# Dynamical regimes of thermally convective emulsions

### v1

@@ II Numerical modeling of emulsions
Figure 1: Characterization of dynamical regimes of thermally convective oil-in-water (O/W) emulsions in the setup of Rayleigh-Bénard thermal convection. Different regimes result from different combinations of the volume fraction of the initially dispersed oil phase $\phi$ and the Rayleigh number $\operatorname{Ra}$ (see text for more details). The characterization is done by considering the presence/absence of thermal convection and the variation in the number of oil droplets $\mathrm{N}_{\mathrm{O}}$ with respect to its initial value $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}$. Panel (a) (grey box): conductive regime, with no variation of $\mathrm{N}_{\mathrm{O}}$. Panel (b) (light-blue box): stable convective regime, with convective plumes and no variation of $\mathrm{N}_{\mathrm{O}}$. Panel (c) (blue box): breakup-dominated convective regime, with convective plumes and an increase of $\mathrm{N}_{\mathrm{O}}$ due to breakup events. Panel (d) (pink box): coalescence-dominated convective regime, with convective plumes and a decrease of $\mathrm{N}_{\mathrm{O}}$ caused by coalescence events. Panel (e) (red box): phase-inverted convective regime, with an initial O/W emulsion that enters a W/O steady convective state. Different colors are used to discriminate between the different regimes. For each pair $(\phi,\operatorname{Ra})$, the corresponding regime in the statistically steady state is shown in Fig. 2.
In this work, we simulate O/W emulsions in a Rayleigh-Bénard setup [98, 99, 100] with a domain of size $L\times H$, with $L\approx 2H$ and wall-to-wall distance $H\approx 20d$, where $d$ is the initial average droplet diameter, which is fixed for all simulations. Two no-slip walls are set at $z=\pm H/2$, and periodic boundary conditions are applied along the $x$ direction. The walls are kept at constant temperatures, $T_{\mathrm{hot}}$ (bottom wall) and $T_{\mathrm{cold}}$ (top wall) (see Fig. 1), corresponding to a temperature jump $\Delta T=T_{\mathrm{hot}}-T_{\mathrm{cold}}$. We follow a dedicated preparation protocol for the emulsions: the oil and water densities are initialized in such a way that an initial number of circular oil droplets, $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}$, corresponding to the desired volume fraction of the dispersed phase, $\phi$, are arranged in a honeycomb-like configuration. The number of initial droplets ranges from $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}=220$ for the most dilute emulsion ($\phi=0.16$) to $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}=700$ for the most packed emulsion ($\phi=0.84$). To create a slightly polydisperse emulsion, we add a small random perturbation to the initial centers-of-mass positions of the droplets and the density field of the continuous phase, and then we leave the emulsion free to relax towards its more energetically favorable configuration [84] (see Fig. 1). We compute the volume fraction of the initially dispersed phase, $\phi$, as the fraction of the domain size initially occupied by the oil phase
$\phi=\frac{A_{\mathrm{O}}}{A_{\mathrm{tot}}},$ (11)
where $A_{\mathrm{O}}$ is the area occupied by the oil phase, computed as in Refs. [71, 72], and $A_{\mathrm{tot}}=L\times H$. After this preparation step, buoyancy forces are added. In a homogeneous fluid, the parameters that control the various dynamical regimes of thermal convection are the Rayleigh number [101, 102, 99, 50, 100]
$\operatorname{Ra}=\frac{\alpha g\Delta TH^{3}}{\nu\kappa},$ (12)
the Prandtl number $\Pr=\nu/\kappa$, and the aspect-ratio $\Gamma=L/H$. In particular, the conduction-to-convection transition occurs as $\operatorname{Ra}$ exceeds a critical threshold [101]. In the case of emulsions, however, as mentioned earlier, interfaces affect the viscous/elastic-visco-plastic characteristics of the system response to the forcing. Therefore, $\operatorname{Ra}$, as written in Eq. (12), does not carry a uniquely determined dynamical information as $\phi$ increases since the effective viscosity acquires a dependency on the local shear rate. For this reason, in the following, we consider the kinematic viscosity of the corresponding homogeneous system in the definition of both $\operatorname{Ra}$ and $\Pr$, and we keep $\operatorname{Ra}$ as a dimensionless measure of the imposed forcing. Notice that the choice of having unitary viscosity and thermal diffusivity ratios, together with imposing no-slip boundary conditions, reduces to four the number of parameters that define the behavior of the emulsion: $\operatorname{Ra}$, $\Pr$, $\Gamma$, and $\phi$. Another relevant parameter is the Weber number, defined as $\operatorname{We}=\rho U^{2}\ell/\Sigma$ (where $U$ is a characteristic flow velocity and $\ell$ a typical droplet-scale length, e.g., the droplet mean radius). However, since we keep the surface tension $\Sigma$ constant, the $\operatorname{We}$ values are related to the $\operatorname{Ra}$ values (we will provide more details about it in the next section). To further simplify the problem, we fix $\Pr=1$ and $\Gamma=2$ in all the simulations and explore the various emerging dynamical regimes in the parameter space spanned by $\operatorname{Ra}$ and $\phi$.
where $\langle\dots\rangle_{x,z}$ denotes the space average. Notice that when emulsions reach their statistically steady states, the Nusselt number in Eq. (13) exhibits fluctuations in time around a mean value [71, 72], hereafter indicated as $\overline{\operatorname{Nu}}$ (see also Sec. III.2):
$\overline{\operatorname{Nu}}=\langle\operatorname{Nu}\rangle_{t}.$
Numerical simulations have been performed on Nvidia V100 and A30 GPUs. We performed about 300 numerical simulations, each typically requiring $\sim 48$ GPU-h of elapsed time. The duration of these simulations in terms of the characteristic free-fall time $t_{FF}\sim\sqrt{H/\alpha g\Delta T}$ depends on the amplitude of the applied buoyancy force. It lays in the range between $t_{\mathrm{FF}}\approx 100$ ($\operatorname{Ra}\approx 4\times 10^{4}$) and $t_{\mathrm{FF}}\approx 4.5\times 10^{3}$ ($\operatorname{Ra}\approx 8\times 10^{6}$). Time-averaged statistics in the statistically steady state (see Sec. III.1) is performed considering intervals of time in the range [$15\div 2000]\,t_{\mathrm{FF}}$, depending on $\operatorname{Ra}$. To analyze droplet sizes (see Sec. IV), we collect data considering all droplets at any time once the statistically steady state is reached.
@@ III.1 Statistically steady states
Previous works [71, 72] showed that the dynamical behavior of concentrated emulsions in a Rayleigh-Bénard setup is strongly affected by both $\phi$ and $\operatorname{Ra}$. However, situations, where the emulsion structure changes during its dynamical evolution via breakup or coalescence events, were not considered, thus focusing only on convective states preserving the number of droplets of the initial condition. In this work, we take a major step forward and complete that analysis by spanning over $\phi\in[0,0.84]$ and $\operatorname{Ra}\in[4\times 10^{3},8\times 10^{6}]$ (correspondingly the range of Weber number is, roughly, $\operatorname{We}\in[10^{-3},10]$) with no discrimination on the structure of the emulsion. For each pair $(\phi,\operatorname{Ra})$, we classify the statistically steady state in terms of the heat flux characteristics (i.e., conduction or convection) and the structural changes in the emulsions, as outlined in Fig. 1. Thus, we can distinguish between states of i) conduction ($\operatorname{Nu}=1$), whereby the emulsion structure is preserved, ii) stable convection ($\operatorname{Nu}>1$), with preserved emulsion structure (i.e., neither droplet breakup nor coalescence is detected), iii) breakup-dominated convection, iv) coalescence-dominated convection, and v) convection in emulsions that undergo a phase inversion. In this scenario, situations explored in Refs. [71, 72] lay in the stable region outlined by $0.1\lesssim\phi\lesssim 0.7$ and $10^{4}\lesssim\operatorname{Ra}\lesssim 6\times 10^{4}$. Indeed, the structural changes take place at relatively high values of $\operatorname{Ra}$, where strong convective flows trigger continuous breakup and coalescence of droplets, eventually leading to a dynamical equilibrium characterized by a mean droplet size $r_{m}$. Notice that, in multiphase turbulent flows, $r_{m}$ corresponds to the so-called Hinze’s scale $r_{H}$ [105, 106], which decreases with the energy dissipation rate, hence with $\operatorname{Ra}$.
@@
Figure 2: Phase diagram reporting a characterization of statistically steady states for different combinations of $\operatorname{Ra}$ and $\phi$ (same symbols/colors as Fig. 1).
@@
For low-to-moderate values of $\phi$, if the initial mean droplet size is larger than $r_{m}$ then breakup overcomes coalescence in the initial transient, due to the low droplet-droplet collision rate, whence we call this regime “breakup-dominated”. Analogously, there exists a range in the pair $(\phi,\operatorname{Ra})$ where the balance in the transient state is broken in favor of coalescence, thus producing “coalescence-dominated” scenarios with fewer larger droplets compared to the initial condition. For emulsions with $\phi$ beyond the jamming point, a sustained convective state is only possible if $\operatorname{Ra}$ is high enough to trigger a phase inversion in the emulsion, as anticipated in Ref. [73]. In Fig. 2, we report the phase diagram of the above-mentioned dynamical regimes. Although we conducted a large number of simulations, they were still not enough to discern whether the transition regions between different regimes are smooth or sharp; therefore, the transition lines drawn in Fig. 2 have just been used to facilitate the localization of the different regimes in the $(\phi,\operatorname{Ra})$ phase diagram. Still, they have not to be understood as sharp transition lines. Note that at a fixed $\phi$ around the jamming point, the system’s behavior as $\operatorname{Ra}$ increases is experiencing a “non-monotonic” sequence: initially, stable convection is observed, which then goes through the coalescence-dominated regime before becoming stable again, eventually leading to phase inversion at higher values of $\operatorname{Ra}$. This non-monotonic behavior arises because of the appearance of the coalescence-dominated regime, which is a precursor of the phase-inverted states, as discussed further in Sec. III.2.
@@
Figure 3: Rheological characterization of emulsions after the preparation step, before buoyancy forces are switched on. We report the shear stress $\sigma$ as a function of the shear rate $\dot{\gamma}$. Panel (a) refers to the lin-lin representation, whereas panel (b) refers to the log-log one. Different symbols/colors correspond to different values of $\phi$ (increasing values from lighter to darker colors). A fit with the Herschel-Bulkley (HB) law (cfr. Eq. (14)) is drawn for $\phi=0.79$ (dashed line).
@@
Figure 5: Characterization of the heat flux in the statistically steady state in terms of the time-averaged Nusselt number $\overline{\operatorname{Nu}}$ as a function of $\operatorname{Ra}$, for different values of $\phi$ (same symbols/colors as Fig. 4). All filled symbols refer to statistically steady states obtained starting from the emulsion at rest and then applying the buoyancy forces. Open crosses in panel (d) refer to statistically steady states obtained starting from an emulsion configuration resulting from a convective statistically steady state and ramping down the intensity of buoyancy forces.
A fundamental aspect is the impact of thermal dynamics on the emulsion rheology. For this purpose, it is necessary to investigate the rheology of the emulsions that reach a statistically steady state compared to the rheology of the emulsions after the preparation step, i.e., before the buoyancy forces are switched on. Hence, we performed simulations with a dedicated Couette rheometer setup without buoyancy forces [84], where walls are moved with constant and opposing velocities along the $x$-axis, ${\bm{u}}_{\mathrm{wall}}=(u_{x}(x,z=\pm H/2,t)=\pm u_{w},0)$. Thus, we impose a shear rate $\dot{\gamma}=2u_{w}/H$ on the emulsion, and then measure the shear stress $\sigma$ after the system has reached the steady state. The corresponding rheological flow curves are reported in Fig. 3. Notice that data for $\phi=0.79$ coincide with the ones reported in the Supplementary Material of Ref. [73], and we stress again the good match with the Herschel–Bulkley (HB) law (dashed black line)
$\sigma=\sigma_{0}+a\dot{\gamma}^{n}\ ,$ (14)
@@
with parameters $\sigma_{0}=2.89\times 10^{-5}$ (i.e., the yield stress), $a=2.2\times 10^{-2}$ (i.e., the consistency index), and $n=0.577$ (i.e., the flux index), thus marking the emergence of a yield stress for high values of $\phi$. We then performed a rheological characterization of the configurations of the statistically steady states. Specifically, we captured the latest emulsion configuration of some representative pairs $(\phi,\operatorname{Ra})$ pinched from the phase diagram, switched off the buoyancy forces, and performed a rheological experiment with the Couette rheometer setup described above. We report the corresponding rheological flow curves in Fig. 4. The flow curve for the breakup-dominated regime (panel (a)) presents an increase of about $20\%$ of the stress at fixed shear rate $\dot{\gamma}$ due to the increase in the number of droplets and, in turn, of the total interfacial energy. Since the breakup-dominated regime comes up only for dilute emulsions, rheological Newtonianity persists. For all other dynamical regimes, i.e., the coalescence-dominated regimes (panels (b) and (c)) and the phase-inverted emulsions (panels (d)-(f)), the activation of the buoyancy force causes a transition from a non-Newtonian to a Newtonian behavior. In the coalescence-dominated regime, the Newtonian behavior is triggered by a decrease in the number of droplets due to coalescence; in the phase inversion regime, phase inversion generates a Newtonian dilute W/O emulsion in marked contrast with the initial non-Newtonian O/W emulsion. To summarize, the resulting emulsions are always Newtonian despite the structural variations induced by the convective states. It is important to remark that this Newtonian behavior cannot be predicted a priori since each pair $(\phi,\operatorname{Ra})$ leads to different emulsion structures regarding the number and size of droplets.
@@
As a further step, we investigated how the aforementioned non-Newtonian-to-Newtonian transition at increasing values of $\operatorname{Ra}$ is reflected in the heat flux properties. We considered the Nusselt number averaged in time over the statistically steady state, $\overline{\operatorname{Nu}}$. Figure  5 shows its relation with $\operatorname{Ra}$ for different values of $\phi$. In all cases (filled symbols), the statistically steady state is reached starting from the emulsion prepared in a static case (with no buoyancy forces) and then applying a given buoyancy amplitude corresponding to the desired $\operatorname{Ra}$. We observe that at increasing $\phi$, the transition to a convective state ($\overline{\operatorname{Nu}}>1$) occurs at higher values of $\operatorname{Ra}$, as expected since the emulsion becomes more viscous at increasing $\phi$. We recall that $\operatorname{Ra}$, which we introduced as an appropriate dimensionless buoyancy force, is defined in terms of the “bare” kinematic viscosity, whereas the ”effective” (dynamically relevant) viscosity grows with $\phi$. Indeed, for what concerns the convective heat flux, emulsions with $\phi$ below jamming behave essentially as a single-phase fluid (but for the shift in the nominal value of the critical $\operatorname{Ra}$), with $\overline{\operatorname{Nu}}$ growing continuously with $\operatorname{Ra}$ above transition. For higher values of $\phi$ (i.e., when the emulsion exhibits yield stress), then the transition to convection is discontinuous, with a jump in $\overline{\operatorname{Nu}}$ witnessing the sudden structural transition from a jammed (conductive) O/W emulsion to a (convective) phase-inverted dilute W/O emulsion whose effective viscosity is reduced to the one dictated by the volume fraction of the water phase. Furthermore, the transition to a phase-inverted state is irreversible, and once the phase inversion is achieved, the system persists in such a Newtonian state. This fact is naturally accompanied by hysteresis in the $\overline{\operatorname{Nu}}$ vs.$\operatorname{Ra}$ curve: starting from an emulsion configuration resulting from a convective state with $\overline{\operatorname{Nu}}>1$ and then ramping down the intensity of buoyancy forces, one observes that $\overline{\operatorname{Nu}}$ goes to zero in a relatively smooth way (panel (d)), echoing the behavior of dilute emulsions. Finally, we note that for $\operatorname{Ra}>5\times 10^{5}$ the curves exhibit a scaling law $\overline{\operatorname{Nu}}\sim\operatorname{Ra}^{\lambda}$, with an exponent close to $\lambda\approx 2/7$, which aligns with findings from studies of thermal convection in simple (single phase) Newtonian fluids [107, 108, 103, 109, 110, 111, 98, 99, 100, 112].

### v2

@@ II.2 Lattice Boltzmann modelling of emulsions
The system domain is resolved with $L\times H=2048\times 1024$ lattice points and we used $\tau=\tau_{h}=1$ in Eqs. (4) and (11). In all simulations, multi-component model parameters are fixed to: $\rho^{\mathrm{max}}_{\mathrm{O,W}}=1.18$, $\rho^{\mathrm{min}}_{\mathrm{O,W}}=0.18$, $\rho_{0}=0.83$, $G_{\mathrm{OW}}=0.405$, $G^{a}_{\mathrm{WW}}=-9.0$, $G^{a}_{\mathrm{OO}}=-8.0$, $G^{r}_{\mathrm{WW}}=8.1$, $G^{r}_{\mathrm{OO}}=7.1$. At the walls, we impose bounce-back boundary conditions for fluid components and Dirichlet boundary conditions for the temperature field. Numerical simulations have been performed on Nvidia V100 and A30 GPUs. We performed about 300 numerical simulations, each typically requiring $\sim 48$ GPU-h of elapsed time. The duration of these simulations in terms of the characteristic free-fall time $t_{FF}\sim\sqrt{H/(\alpha g\Delta T)}$ depends on the amplitude of the applied buoyancy force. It lays in the range between $t_{\mathrm{FF}}\approx 100$ ($\operatorname{Ra}\approx 4\times 10^{4}$) and $t_{\mathrm{FF}}\approx 4.5\times 10^{3}$ ($\operatorname{Ra}\approx 8\times 10^{6}$). Time-averaged statistics in the statistically steady state (see Sec. III.1) is performed considering intervals of time in the range [$15\div 2000]\,t_{\mathrm{FF}}$, depending on $\operatorname{Ra}$. To analyze droplet sizes (see Sec. IV), we collect data considering all droplets at any time once the statistically steady state is reached.
@@
Figure 1: Characterization of dynamical regimes of thermally convective oil-in-water (O/W) emulsions in the setup of Rayleigh-Bénard thermal convection. Different regimes result from different combinations of the volume fraction of the initially dispersed oil phase $\phi$ and the Rayleigh number $\operatorname{Ra}$ (see text for more details). The characterization is done by considering the presence/absence of thermal convection and the variation in the number of oil droplets $\mathrm{N}_{\mathrm{O}}$ with respect to its initial value $\mathrm{N}^{\mathrm{init}}_{\mathrm{O}}$. Panel (a) (grey box): conductive regime, with no variation of $\mathrm{N}_{\mathrm{O}}$. Panel (b) (light-blue box): stable convective regime, with convective plumes and no variation of $\mathrm{N}_{\mathrm{O}}$. Panel (c) (blue box): breakup-dominated convective regime, with convective plumes and an increase of $\mathrm{N}_{\mathrm{O}}$ due to breakup events. Panel (d) (pink box): coalescence-dominated convective regime, with convective plumes and a decrease of $\mathrm{N}_{\mathrm{O}}$ caused by coalescence events. Panel (e) (red box): phase-inverted convective regime, with an initial O/W emulsion that enters a W/O steady convective state. Different colors are used to discriminate between the different regimes. For each pair $(\phi,\operatorname{Ra})$, the corresponding regime in the statistically steady state is shown in Fig. 2. Movies corresponding to panels (b)-(e) can be found in the Supplementary Material.
@@ III.1 Statistically steady states
Previous works [71, 72] showed that the dynamical behavior of concentrated emulsions in a Rayleigh-Bénard setup is strongly affected by both $\phi$ and $\operatorname{Ra}$. However, situations where the emulsion structure changes during its dynamical evolution via breakup or coalescence events have not been considered, thus focusing only on convective states preserving the number of droplets of the initial condition. In this work, we take a major step forward and complete that analysis by spanning over $\phi\in[0,0.84]$ and $\operatorname{Ra}\in[4\times 10^{3},8\times 10^{6}]$ (correspondingly the range of Weber number is, roughly, $\operatorname{We}\in[10^{-3},10]$) with no discrimination on the structure of the emulsion. For each pair $(\phi,\operatorname{Ra})$, we classify the statistically steady state in terms of the heat flux characteristics (i.e., conduction or convection) and the structural changes in the emulsions, as outlined in Fig. 1. Thus, we can distinguish between states of i) conduction ($\operatorname{Nu}=1$), whereby the emulsion structure is preserved, ii) stable convection ($\operatorname{Nu}>1$), with preserved emulsion structure (i.e., neither droplet breakup nor coalescence is detected), iii) breakup-dominated convection, iv) coalescence-dominated convection, and v) convection in emulsions that undergo a phase inversion. In this scenario, situations explored in Refs. [71, 72] lay in the stable region outlined by $0.1\lesssim\phi\lesssim 0.7$ and $10^{4}\lesssim\operatorname{Ra}\lesssim 6\times 10^{4}$. Indeed, the structural changes take place at relatively high values of $\operatorname{Ra}$, where strong convective flows trigger continuous breakup and coalescence of droplets, eventually leading to a dynamical equilibrium characterized by a mean droplet size $r_{m}$. Notice that, in multiphase turbulent flows, $r_{m}$ corresponds to the so-called Hinze’s scale $r_{H}$ [110, 111], which decreases with the energy dissipation rate, hence with $\operatorname{Ra}$.
@@
Figure 2: Phase diagram reporting a characterization of statistically steady states for different combinations of $\operatorname{Ra}$ and $\phi$ (same colors as Fig. 1).
@@
For low-to-moderate values of $\phi$, if the initial mean droplet size is larger than $r_{m}$ then breakup overcomes coalescence in the initial transient, due to the low droplet-droplet collision rate, whence we call this regime “breakup-dominated”. Analogously, there exists a range in the pair $(\phi,\operatorname{Ra})$ where the balance in the transient state is broken in favor of coalescence, thus producing “coalescence-dominated” scenarios with fewer larger droplets compared to the initial condition. For emulsions with $\phi$ beyond the jamming point, a sustained convective state is only possible if $\operatorname{Ra}$ is high enough to trigger a phase inversion in the emulsion, as anticipated in Ref. [73]. In Fig. 2, we report the phase diagram of the above-mentioned dynamical regimes. Although we conducted a large number of simulations, they were still not enough to discern whether the transition regions between different regimes are smooth or sharp; therefore, the transition lines drawn in Fig. 2 have just been used to facilitate the localization of the different regimes in the $(\phi,\operatorname{Ra})$ phase diagram. Still, they have not to be understood as sharp transition lines. Note that at a fixed value of $\phi$ around the jamming point, the system’s behavior as $\operatorname{Ra}$ increases is experiencing a “non-monotonic” sequence: initially, stable convection is observed, which then goes through the coalescence-dominated regime before becoming stable again, eventually leading to phase inversion at higher values of $\operatorname{Ra}$. This non-monotonic behavior arises because of the appearance of the coalescence-dominated regime, which is a precursor of the phase-inverted states, as discussed further in Sec. III.2.
A fundamental aspect is the impact of thermal dynamics on the emulsion rheology. For this purpose, it is necessary to investigate the rheology of the emulsions that reach a statistically steady state compared to the rheology of the emulsions after the preparation step, i.e., before the buoyancy forces are switched on. Hence, we performed simulations with a standard Couette rheometer setup without buoyancy forces [88], where walls are moved with constant and opposing velocities along the $x$-axis, ${\bm{u}}_{\mathrm{wall}}=(u_{x}(x,z=\pm H/2,t)=\pm u_{w},0)$. Thus, we impose a shear rate $\dot{\gamma}=2u_{w}/H$ on the emulsion, and then we measure the shear stress $\sigma$ after the system has reached the statistically steady state. The corresponding rheological flow curves are reported in Fig. 3. Notice that data for $\phi=0.79$ coincide with the ones reported in the Supplementary Material of Ref. [73], and we stress again the good match with the Herschel–Bulkley (HB) law (dashed black line)
$\sigma=\sigma_{0}+a\dot{\gamma}^{n}\ ,$ (14)
with parameters $\sigma_{0}=2.89\times 10^{-5}$ (i.e., the yield stress), $a=2.2\times 10^{-2}$ (i.e., the consistency index), and $n=0.577$ (i.e., the flux index), thus marking the emergence of a yield stress for high values of $\phi$. We then performed a rheological characterization of the configurations of the statistically steady states. Specifically, we captured the latest emulsion configuration of some representative pairs $(\phi,\operatorname{Ra})$ pinched from the phase diagram, switched off the buoyancy forces, and performed a rheological experiment with the Couette rheometer setup described above.
@@
Figure 3: Rheological characterization of emulsions after the preparation step, before buoyancy forces are switched on. We report the shear stress $\sigma$ as a function of the shear rate $\dot{\gamma}$. Panel (a) refers to the lin-lin representation, whereas panel (b) refers to the log-log one. Different symbols/colors correspond to different values of $\phi$ (increasing values from lighter to darker colors). A fit with the Herschel-Bulkley (HB) law (cfr. Eq. (14)) is drawn for $\phi=0.79$ (dashed black line).
@@
Figure 5: Characterization of the heat flux in the statistically steady state in terms of the Nusselt number averaged in time over the statistically steady state, $\overline{\operatorname{Nu}}$, as a function of $\operatorname{Ra}$, for different values of $\phi$ (same symbols/colors as Fig. 3). All filled symbols refer to statistically steady states obtained starting from the emulsion at rest and then applying the buoyancy forces. Open crosses in panel (d) refer to statistically steady states obtained starting from an emulsion configuration resulting from a convective statistically steady state and ramping down the intensity of buoyancy forces.
@@
We report the corresponding rheological flow curves in Fig. 4. The flow curve for the breakup-dominated regime (panel (a)) presents an increase of about $20\%$ of the stress at fixed shear rate $\dot{\gamma}$ due to the increase in the number of droplets and, in turn, of the total interfacial energy. Since the breakup-dominated regime comes up only for dilute emulsions, rheological Newtonianity persists. For all other dynamical regimes, i.e., the coalescence-dominated regimes (panels (b) and (c)) and the phase-inverted emulsions (panels (d)-(f)), the activation of the buoyancy force causes a transition from a non-Newtonian to a Newtonian behavior. In the coalescence-dominated regime, the Newtonian behavior is triggered by a decrease in the number of droplets due to coalescence; in the phase inversion regime, phase inversion generates a Newtonian dilute W/O emulsion in marked contrast with the initial non-Newtonian O/W emulsion. To summarize, the resulting emulsions are always Newtonian despite the structural variations induced by the convective states. It is important to remark that this Newtonian behavior cannot be predicted a priori since each pair $(\phi,\operatorname{Ra})$ leads to different emulsion structures regarding the number and size of droplets.
@@
As a further step, we investigated how the aforementioned non-Newtonian-to-Newtonian transition at increasing values of $\operatorname{Ra}$ is reflected in the heat flux properties. We considered the Nusselt number averaged in time over the statistically steady state, $\overline{\operatorname{Nu}}$. Figure  5 shows its relation with $\operatorname{Ra}$ for different values of $\phi$. In all cases (filled symbols), the statistically steady state is reached starting from the emulsion prepared in a static case (with no buoyancy forces) and then applying a given buoyancy amplitude corresponding to the desired $\operatorname{Ra}$. We observe that at increasing $\phi$, the transition to a convective state ($\overline{\operatorname{Nu}}>1$) occurs at higher values of $\operatorname{Ra}$, as expected since the emulsion becomes more viscous at increasing $\phi$. We recall that $\operatorname{Ra}$, which we introduced as an appropriate dimensionless buoyancy force, is defined in terms of the “bare” kinematic viscosity, whereas the “effective” (dynamically relevant) viscosity grows with $\phi$. Indeed, for what concerns the convective heat flux, emulsions with $\phi$ below jamming behave essentially as a single-phase fluid (but for the shift in the nominal value of the critical $\operatorname{Ra}$), with $\overline{\operatorname{Nu}}$ growing continuously with $\operatorname{Ra}$ above transition. For higher values of $\phi$ (i.e., when the emulsion exhibits yield stress), then the transition to convection is discontinuous, with a jump in $\overline{\operatorname{Nu}}$ witnessing the sudden structural transition from a jammed (conductive) O/W emulsion to a (convective) phase-inverted diluted W/O emulsion whose effective viscosity is reduced to the one dictated by the volume fraction of the water phase. Furthermore, the transition to a phase-inverted state is irreversible, and once the phase inversion is achieved, the system persists in such a Newtonian state. This fact is naturally accompanied by hysteresis in the $\overline{\operatorname{Nu}}$ vs.$\operatorname{Ra}$ curve: starting from an emulsion configuration resulting from a convective state with $\overline{\operatorname{Nu}}>1$ and then ramping down the intensity of buoyancy forces, one observes that $\overline{\operatorname{Nu}}$ goes to zero in a relatively smooth way (panel (d)), echoing the behavior of dilute emulsions. Finally, we note that for $\operatorname{Ra}>5\times 10^{5}$ the curves exhibit a scaling law $\overline{\operatorname{Nu}}\sim\operatorname{Ra}^{\lambda}$, with an exponent close to $\lambda\approx 2/7$, which aligns with findings from studies of thermal convection in simple (single phase) Newtonian fluids [112, 113, 92, 114, 115, 116, 84, 85, 86, 117].
III.2 Transient dynamics and heat flux fluctuations
We complement the statistically steady state characterization of Sec. III.1 with insights into the transient dynamics and heat flux fluctuations. At first, to have a reference case to compare with, we simulated the dynamics of a Newtonian homogeneous fluid under thermal convection in the Rayleigh-Bénard setup. In Fig. 6, we report the time evolution of the Nusselt number, $\operatorname{Nu}$, for different values of $\operatorname{Ra}$. As expected [89], for values of $\operatorname{Ra}$ above a critical value, we observe stable convection in a wide range of values of $\operatorname{Ra}$, roughly between $\operatorname{Ra}\approx 2\times 10^{3}$ and $\operatorname{Ra}\approx 10^{6}$. In this range, the transient dynamics of $\operatorname{Nu}$ shows oscillations whose amplitude decays exponentially over time [118]. Above this upper value, oscillations during transient dynamics are not damped anymore but persist with a periodic pattern, a scenario that is symptomatic of the progressive approach to turbulence [119]. We then select some representative pairs $(\phi,\operatorname{Ra})$ and report the corresponding time evolution of $\operatorname{Nu}$ in Fig. 7. In particular, we consider a dilute emulsion showing a breakup-dominated regime ($\phi=0.16$, panel (a)); a semi-dilute emulsion still showing a breakup-dominated regime ($\phi=0.39$, panel (b)); a concentrated emulsion, just below the jamming point ($\phi=0.68$, panel (c)); a jammed emulsion, just above the jamming point ($\phi=0.79$, panel (d)); a highly packed, jammed emulsion ($\phi=0.84$, panel (e)). Regular oscillations, which are peculiar to the Newtonian homogeneous fluid (see Fig. 6), tend to acquire a stochastic component when $\phi>0$, both in the transient and in the statistically steady state. This randomness is rooted in the presence of finite-sized droplets, which collide while transported by the flow. For fixed $\operatorname{Ra}$, fluctuations in $\operatorname{Nu}$ increase at increasing $\phi$: this is particularly evident when $\operatorname{Ra}\approx 4\times 10^{6}$ and we approach conditions triggering the phase inversion (see panels (a)-(c)). This behavior is similar to what has been observed in a recent study on a Taylor-Couette isothermal system [83], where increasing drag fluctuations are observed while approaching the catastrophic phase inversion. Interestingly, signatures of a beat pattern emerge for $\operatorname{Ra}\approx 4\times 10^{6}$ and are more pronounced for the intermediate values of $\phi$. The combination ($\phi=0.79$, $\operatorname{Ra}\approx 4\times 10^{5}$) reported in panel (d) coincides with the regime explored in Ref. [73], with intermittency characterizing the transient dynamics, where long conductive periods with $\operatorname{Nu}\approx 1$ alternate with intermittent heat bursts where $\operatorname{Nu}\gg 1$, overall resulting in a very long transient dynamics (much longer than the time interval shown in Fig. 7).